# Influence of adversarial training on super-resolution turbulence reconstruction

### v4

@@ Acknowledgments
The research leading to these results has received funding from the European Union’s Horizon 2020 research and innovation program under the Center of Excellence in Combustion (CoEC) project, grant agreement no. 952181, and from the German Federal Ministry of Education and Research (BMBF) and the state of North Rhine-Westphalia as part of the NHR funding. The authors gratefully acknowledge the computing resources from the DEEP-EST project, which received funding from the European Union’s Horizon 2020 research and innovation program under grant agreement no. 754304 and the computing time granted by the NHR4CES Resource Allocation Board on the high-performance computer CLAIX at the NHR Center RWTH Aachen University (project rwth0733). We thank Dr. R. Sedona and Dr. G. Cavallaro for their support in the porting of the application to DEEP-EST. The authors thank Mr. S. Sakhare, Mr. F. Fröde, and Dr. P. Petkov for their exceptional support and contributions to this research project.

### v5

@@ Acknowledgments
The research leading to these results has received funding from the European Union’s Horizon 2020 research and innovation program under the Center of Excellence in Combustion (CoEC) project, Grant Agreement No. 952181, and from the German Federal Ministry of Education and Research (BMBF) and the state of North Rhine-Westphalia as part of the NHR funding. The authors gratefully acknowledge the computing resources from the DEEP-EST project, which received funding from the European Union’s Horizon 2020 research and innovation program under Grant Agreement no. 754304 and the computing time granted by the NHR4CES Resource Allocation Board on the high-performance computer CLAIX at the NHR Center RWTH Aachen University (project rwth0733). We thank Dr. R. Sedona and Dr. G. Cavallaro for their support in the porting of the application to DEEP-EST. The authors thank S. Sakhare, F. Fröde, and Dr. P. Petkov for their exceptional support and contributions to this research project.